OpenStudy (anonymous):

Determine the standard form of the equation of the line that passes through (2, 7) and (-6,0)

OpenStudy (anonymous):

Walk Me Through Please !

OpenStudy (arhin):

ayt

OpenStudy (arhin):

find the slope first

OpenStudy (anonymous):

m = -7/4 @Arhin

OpenStudy (arhin):

choose one of the points

OpenStudy (arhin):

and fix it in the formula y - \[y-y _{1}=\frac{ dy }{ dx }(x-x _{1})\]

OpenStudy (anonymous):

um so lets use (2,7) im confused just a little

OpenStudy (arhin):

where x1 and y1 are the coordinates of the point you chose

OpenStudy (arhin):

\[y-7=\frac{ -7 }{ 4 }(x-2)\]

OpenStudy (arhin):

simplify your answer and write it in the form \[ax+by+c=0\]

OpenStudy (anonymous):

@Arhin ok so I know its gonna be -7x

OpenStudy (arhin):

\[-7x -4y-14=0 \]

OpenStudy (arhin):

or \[7x+4y+14=0\]

OpenStudy (anonymous):

Mhm these are not one of my choices given I have -7x + 8y = 42 -7x - 8y = 42 -7x + 8y = -42 8x + 7y = 42 @Arhin

OpenStudy (arhin):

yeah sorry the slope is rather 7/8

OpenStudy (arhin):

you will get A as your answer